OpenStudy (anonymous):

If x^4 = 1,296 what does x equal?

OpenStudy (anonymous):

What is \[ 6^4? \]

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

\[ 6^4= 1296 \]

OpenStudy (anonymous):

If you needed proof 6*6 = 36 36*6 = 216 216*6 = 1296